OpenStudy (anonymous):

every Monday, Rajiv receives fifty percent off enchilada dinners that regularly cost thirty-two dollars. Which of the following shows the amount Rajiv pays for the dinners if he tips the waiter seventeen percent of the regular price for the dinners? Assume no sales tax is added to his bill. sixteen dollars and seventeen cents twenty-one dollars and forty-four cents thirty-two dollars and seventeen cents thirty-seven dollars and forty-four cents

OpenStudy (help!!!!):

.17*32=??

OpenStudy (anonymous):

?

OpenStudy (help!!!!):

What is .17*32?

OpenStudy (anonymous):

544

OpenStudy (help!!!!):

are you serious?

OpenStudy (help!!!!):

It's 5.44

OpenStudy (anonymous):

same thing

OpenStudy (help!!!!):

no they are VERY different

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

26.56?

OpenStudy (help!!!!):

wait asec

OpenStudy (help!!!!):

you actually add 5.44+32

OpenStudy (help!!!!):

SO what we did was find 17% of $32 and we added it to 32

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

Tori's twelve-ounce glass of soymilk has 9 grams of protein. Brady's sixteen-ounce glass has twelve grams of protein. What is the ratio of the protein in Brady's drink to the protein in Tori's drink? one colon one three colon four four colon three sixteen colon four

OpenStudy (texaschic101):

its 17% of regular price...so 17% of 32 = 0.17 * 32 = $5.44 ======================== protein in Brandy's = 9 protein in Tori's = 12 9:12 is the same as 9/12, which reduces to 3/4, which is the same as 3:4 (your answer) sometimes they will throw in extra info into the problem...info that is not needed to do the problem

OpenStudy (texaschic101):

any questions ?

OpenStudy (texaschic101):

wait...I read the first one wrong...let me check it again

OpenStudy (anonymous):

wow really

OpenStudy (texaschic101):

He tips waitor 5.44.....he usually pays 1/2 of 32 which is 16.....16 + 5.44 = $21.44

OpenStudy (anonymous):

thats not an answere?

OpenStudy (texaschic101):

second answer choice is $21.44

OpenStudy (anonymous):

@Kaylee_td

OpenStudy (texaschic101):

your second question...answer is 3:4

OpenStudy (texaschic101):

do you have any questions about any of them ?

OpenStudy (anonymous):

what about this Match each actual value with the predicted value that results in a fifteen percent error. Match Term Definition Actual: eighty A) Predicted: ninety−two Actual: one hundred twenty B) Predicted: one hundred two Actual: one hundred C) Predicted: one hundred nineteen Actual: one hundred forty D) Predicted: eighty−five

OpenStudy (texaschic101):

It does not say if the error is positive or negative 0.15(80) + 80 = 12 + 80 = 92 (15% increase) -- A 120 - 0.15(120) = 120 - 18 = 102 (15% decrease) -- B 100 = 0.15(100) = 100 - 15 = 85 (15% decrease) -- D 140 = 0.15(140) = 140 - 21 = 119 (15% decrease) -- C sorry it took so long...had to figure out how to do it